# Meta-Reasoner: Dynamic Guidance for Optimized Inference-time Reasoning in Large Language Models

### v3

@@ 4.3 Meta-reasoner Strategy Generation
In the next step, the meta-reasoner evaluates the progress report $P_{t}$ and selects an appropriate strategy $G_{t}$ for LLM reasoning (the complete procedure is detailed in Algorithm 1). We formulate the generation of strategy as a contextual MAB problem and consider two settings: (1) a fixed-strategy formulation, where the meta-reasoner selects from a predefined set of strategies using a contextual bandit algorithm; and (2) an advanced setting, in which the meta-reasoner itself is an LLM-based agent capable of introducing or refining strategies dynamically. In both cases, the meta-reasoner employs the partial-feedback principle of MABs to adaptively select strategies based on a reward function that evaluates the quality of reasoning progress after applying the chosen strategy. We demonstrate the contextual bandit pair (i.e., diagnosis of the progress report (i.e., context) and the corresponding strategy (i.e., bandit) in Table 1.
Progress Evaluation.
@@
A central objective of our evaluation mechanism is to quantify how effectively the model’s current reasoning advances toward the task goal (e.g., solving a complex problem), while also monitoring computational cost to promote efficiency. Concretely, we implement a reward function that tracks both solution progress (e.g., partial correctness, adherence to constraints) and resource usage (e.g., number of reasoning steps). This evaluator can be any suitable mechanism, including LLM-based verification or external scoring scripts. In our implementation, we leverage an LLM as the evaluator (with prompts referenced in Figures 47) to assess reasoning progress. The evaluator outputs a cumulative score, which is then used to update the MAB algorithm.
Fixed Contextual Bandit.
@@
In the basic version of our framework, the meta-reasoner is modeled as a single contextual bandit that selects from a fixed, finite set of $K$ strategies. These strategies include instructions such as “continue and provide specific suggestions”, “restart from scratch”, “backtrack to the point where the error occurred”, or “propose alternative methods or perspectives to consider”, as detailed in Table 1 (rows without marked). At each round, the LLM produces a progress report summarizing its partial reasoning, the meta-reasoner transforms this progress report into a feature vector $x_{t}$ using a language model and applies a contextual bandit algorithm (e.g., LinUCB Li et al. (2012)) to select the best next strategy $a_{t}$. The LLM then executes that strategy and we collect the reward $r_{t}$ for $a_{t}$ based on the reward function. Through iterative MAB algorithm updating, the MAB algorithm learns to select appropriate strategies conditioned on the recent progress report.
Dynamic Contextual Bandit.
@@
The fixed-arm formulation assumes a static set of strategies. In practice, the meta-reasoner may itself be an LLM capable of inventing new strategies over time. To accommodate dynamic strategies, we allow the meta-reasoner to propose or refine new strategies at round $t$, which generates an expanding collection of strategies, $G_{1}\subseteq\cdots\subseteq G_{t}$. Each newly proposed strategy becomes an arm in the contextual bandit. To encourage at least some exploration on this new arm, we initialize each arm with a blank or weak prior in the bandit’s parameters. We further analyze the stability of these new generated dynamic contextual bandits in Appendix §C.
@@
By explicitly separating low-level content generation (handled by the LLM) from high-level strategy decisions (governed by the meta-reasoner’s bandit), the system can effectively avoid getting stuck or wasting excessive resources on poor solution paths. In domains where a predefined set of strategies is sufficient, the fixed-arm formulation can simplify the method deployment. While in more open-ended domains where novel tactics may emerge, dynamic-arm extensions give meta-reasoner more flexibility to evolve.
5 Experiments
@@
In this section, we first introduce the experiment settings including datasets, backbone models and training details of MAB. We then present the main results of Meta-Reasoner with analysis regarding efficiency, rewards accumulation, and qualitative assessment of meta-reasoner output. We present the detailed baselines in Appendix A.
5.1 Experiments Setup
@@
We test our method on several challenging datasets that demand complex reasoning and often involve lengthy thinking processes for the correct solutions. These includes (1) 24-point game Yao et al. (2023); (2) college-level scientific problem from SciBench Wang et al. (2024) and (3) math questions based on theorems from TheoremQA Chen et al. (2023). For SciBench, we focus only on the math-related subsets (i.e., diff, stat, and calc). Detailed explanations for each subset can be found in Wang et al. (2024). For TheormQA, we only consider the math subset that involves logical reasoning.
@@
We collect the training data for each task: (1) for the 24-point game, we random sample 50 queries from 4nums.com specifically excluding problems in ranks 901-1000 which were reserved for testing; (2) for TheoremQA, we randomly sample 30 mathematical reasoning queries from the dataset; (3) for SciBench, we randomly sample 30 queries from differential subsets including diff, stat, and calc from the entire dataset. These samples were used to iteratively update the LinUCB parameters for both fixed ($K$=3 or $K$=5 strategies) and dynamic strategy settings.
@@
We configure the training process using deterministic generation ($n$=1, Top_k=1, temperature=0) with specific max_token limits for CoT generation (512), meta-reasoner feedback (256), progress reports (512), and reward model outputs (4). The LinUCB exploration parameter was set to $c$=0.2. Experiments were ran for a maximum of $T$=30 iterations for the 24-point game, scibench, and $T$=100 for TheoremQA, with MAB parameters updated after each iteration based on a reward function that weighted objective completion (40%), progress quality (30%), efficiency (15%), and strategy alignment (15%). The reward function prompt is detailed in Figure 4-7.
@@
[…]
GPT-4o-mini + MACM Lei et al. (2024) 54.78 67.13 65.77
GPT-4o + MACM Lei et al. (2024) 61.42 78.32 76.72
@@
Table 2: Accuracy (%) comparison of different methods on the math-related subset of the SciBench dataset. Each column refers to the problem subset defined in Wang et al. (2024).
Method Accuracy (%)
GPT-4o-mini + CoT Yao et al. (2023) 4
GPT-4o-mini + SC-CoT Yao et al. (2023) 9
GPT-4o-mini + IO (best of 100) Yao et al. (2023) 33
GPT-4o-mini + CoT (best of 100) Yao et al. (2023) 49
Gemini-Exp-1206 + IO (best of 100) Yao et al. (2023) 38
Gemini-Exp-1206 + CoT (best of 100) Yao et al. (2023) 60
[…]
GPT-4o-mini + ToT ($b=1$Yao et al. (2023) 32
GPT-4o-mini+ ToT ($b=5$Yao et al. (2023) 65
GPT-4o-mini + Reflexion Shinn et al. (2024) 53
GPT-4o-mini + MACM Lei et al. (2024) 80
GPT-4o-mini + Meta-Reasoner (our work) 89
GPT-4o + Meta-Reasoner (our work) 92
Gemini-Exp-1206 + Meta-Reasoner (our work) 94
[…]
o1-mini + IO 89
o1-preview + IO 93
Table 3: Accuracy(%) comparison of different prompting methods on 24-points game. $b$: Search breadth.
Backbone Models.
We consider both LLMs and the recent Large Reasoning Models (LRMs) for our experiments. For the LLMs, we consider the closed-source models like gpt-4o, gpt-4o-mini (between Nov 2025 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-3.1-8B-instruct from Meta, phi-4 from Microsoft and gemini-experimental-1206 from Google. For the LRMs, we consider the closed-source models like o1, o1-mini (In case we cannot break down the generation of o1 models through APIs, we cannot properly inject our meta-reasoner with o1-series models; we only provide the IO results for references). For the feature extraction mentioned in §4.3, we use text-embedding-3-small from OpenAI as the embedding model.
To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models. We use the API service from OpenAI [^2] https://openai.com/ and OpenRouter [^3] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
5.2 Main Results
We compare the accuracy of different prompting methods across different backbone models on SciBench (as shown in Table 2), 24-points game (as shown in Table 3) and TheoremQA (as shown in Table 4). We found that basic prompting strategies, such as CoT and SC-CoT, show limited effectiveness, achieving only 4% and 9% accuracy on 24-point games, respectively. Incorporating IO strategy with “Best of 100” samples improves accuracy to 33%, but it remains far behind advanced methods. Strategies like ToT illustrate the importance of exploring broader reasoning paths, with accuracy increasing from 45% to 74% as the search breadth expands from 1 to 5. Advanced iterative methods, such as Reflexion (53%) and MACM (80%), further demonstrate the value of refined reasoning frameworks. Our proposed Meta-Reasoner outperforms these approaches, achieving 89% accuracy with GPT-4o-mini and 92% with GPT-4o, showcasing its ability to dynamically guide reasoning, correct errors, and focus resources effectively. Compared to specialized models like o1-mini, our method equipped with much cheaper and generalized models like GPT-4o-mini delivers comparable performance, demonstrating its adaptability and scalability. Overall, the Meta-Reasoner framework provides a compatible approach to improving reasoning-intensive tasks, combining high accuracy with dynamic and efficient problem-solving strategies. The results on SciBench and TheoremQA also demonstrate similar findings and show that Meta-Reasoner generally achieves better performance compared to the baselines and the results are consistent across different models.
5.3 Ablation Study
Method Accuracy (%)
GPT-4o-mini + CoT 39.46
Gemini-Exp-1206 + CoT 43.12
[…]
GPT-4o-mini + Reflexion Shinn et al. (2024) 74.32
GPT-4 Turbo + MACM Lei et al. (2024) 79.41
GPT-4o-mini + Meta-Reasoner (our work) 84.13
Gemini-Exp-1206 + Meta-Reasoner (our work) 86.32
Table 4: Accuracy(%) comparison of different prompting methods on TheoremQA Chen et al. (2023).
Model Variant Game-of-24(%) TheoremQA(%)
@@ 5.3 Ablation Study
Table 5: Ablation study of Meta-Reasoner. Direct arm selection refers to prompting LLM to directly select a strategy based on recent progress report.

### v4

@@ 4.3. Meta-reasoner Strategy Generation
In the next step, the meta-reasoner evaluates the progress report $P_{t}$ and selects an appropriate strategy $G_{t}$ for LLM reasoning (the complete procedure is detailed in Algorithm 1). We formulate the generation of strategy as a CMAB problem (as defined in §3) and explore two settings: (1) a fixed-strategy formulation, where the meta-reasoner selects from a predefined set of strategies using a contextual bandit algorithm; and (2) an advanced setting, where the meta-reasoner, implemented as an LLM-based agent, dynamically generates or refines strategies. In both settings, the meta-reasoner leverages the partial-feedback mechanism of MABs to adaptively select strategies based on a reward function that evaluates the quality of reasoning progress after applying the chosen strategy $G_{t}$. We demonstrate the contextual bandit pair (diagnosis of the progress report (i.e., context) and the corresponding strategy (i.e., bandit) in Table 1.
Algorithm 1 Meta-Reasoning with Contextual Multi-Armed Bandits
0:  LLM $M$, Bandits $\mathcal{B}$, Initial strategy set $\mathcal{A}_{1}$, Maximum rounds $T$
0:  Final answer $A_{\mathrm{final}}$
1:  $C_{0}\leftarrow\emptyset$;  $\mathcal{B}.\mathrm{Initialize}(\mathcal{A}_{1})$
2:  $G_{0}\leftarrow$ default strategy
3:  for $t=1$ to $T$ do
4:   if $t>1$ then
5:    $P_{t-1}\leftarrow f(C_{t-1})$
6:    $x_{t-1}\leftarrow\mathrm{FeatureExtract}(P_{t-1})$
7:    (Optional): $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t-1}\cup\{\text{new strategies}\}$
8:    $a_{t-1}\leftarrow\arg\max_{a\in\mathcal{A}_{t}}\mathrm{Score}_{\mathcal{B}}(x_{t-1},a)$
9:    $G_{t}\leftarrow a_{t-1}$
10:   else
11:    $G_{t}\leftarrow G_{0}$
12:   end if
13:   $s_{t}\leftarrow M(C_{t-1},G_{t})$
14:   $C_{t}\leftarrow C_{t-1}\cup\{s_{t}\}$
15:   $r_{t}\leftarrow\mathrm{ComputeReward}(C_{t})$
16:   if $t>1$ then
17:    $\mathcal{B}.\mathrm{Update}(x_{t-1},a_{t-1},r_{t})$
18:   end if
19:   if termination condition met then
20:    break
21:   end if
22:  end for
23:  $A_{\mathrm{final}}\leftarrow\mathrm{ExtractAnswer}(C_{t})$
24:  return  $A_{\mathrm{final}}$
4.3.1. Reward Modeling for Progress Report.
@@
The primary goal of our evaluation mechanism is to quantify how effectively the model’s current reasoning advances toward the task objective (e.g., solving a complex problem), while also monitoring computational cost to promote efficiency. We deine a reward function $R:R:P_{t}\times G_{t}\to\mathbb{R}$ that integrates two key components: (1) solution progress $S_{p}$, measuring correctness and adherence to query’s constraints (e.g., $S_{p}=w_{1}\cdot C_{c}+w_{2}\cdot C_{a}$, where $C_{c}$ is correctness, $C_{a}$ is adherence, and $w_{1},w_{2}$ are weights); and (2) resource usage $R_{u}$, reflecting computational cost (e.g., $R_{u}=-\alpha\cdot N_{s}$, where $N_{s}$ is the number of reasoning steps and $\alpha$ is a cost coefficient). The total reward is computed as $R(G_{t},P_{t})=\beta\cdot S_{p}+(1-\beta)\cdot R_{u}$, with $\beta\in[0,1]$ balancing the trade-off. This evaluation can be implemented via an LLM-based assessor (prompts detailed in Figures 69) or external scoring scripts, producing a cumulative score to update the CMAB algorithm. For our experiments, we use the same model as both CoT generation and reward modeling.
4.3.2. Fixed Contextual Bandit.
@@
In the basic version of our framework, the meta-reasoner is modeled as a single contextual bandit that selects from a fixed, finite set of $K$ strategies. These strategies include instructions such as “continue and provide specific suggestions”, “restart from scratch”, “backtrack to the point where the error occurred”, or “propose alternative methods or perspectives to consider”, as detailed in Table 1. At each round, the LLM produces a progress report $P_{t}$ summarizing its recent several reasoning trajectory, the meta-reasoner then transforms this progress report into a feature vector $x_{t}$ using a language model and applies a contextual bandit algorithm (e.g., LinUCB (Li et al., 2010)) to select the best next strategy $a_{t}$. The LLM then executes that strategy and we collect the reward $r_{t}$ for $a_{t}$ based on the reward function. Through iterative MAB algorithm updating, the MAB algorithm learns to select appropriate strategies conditioned on the recent progress report.
4.3.3. Dynamic Contextual Bandit.
@@
The fixed-arm formulation assumes a static set of strategies $G_{0}$. In practice, the meta-reasoner may itself be an LLM capable of inventing new strategies over time. To accommodate dynamic strategies, we allow the meta-reasoner to propose or refine new strategies at round $t$, which generates an expanding collection of strategies, $G_{1}\subseteq G_{2}\subseteq\cdots\subseteq G_{t}$. Each newly introduced strategy becomes an additional arm in the contextual multi-armed bandit framework. To encourage at least some exploration on this new arm, we initialize their bandit parameters with a neutral prior, such as $Q(a)=0$ for arm $a$ to avoid strong biases. We further analyze the stability of these new generated dynamic contextual bandits in §5.4.5, ensuring that the system avoids incorporating random or ineffective strategies by grounding new arms in contextual relevance and iterative feedback.
@@
By explicitly separating low-level content generation (handled by the LLM) from high-level strategy decisions (governed by the meta-reasoner’s bandit), the system can effectively avoid getting stuck or wasting excessive resources on poor reasoning paths. In domains where a predefined set of strategies is sufficient, the fixed-arm formulation can simplify the method deployment. While in more open-ended domains where novel tactics may emerge, dynamic-arm extensions give meta-reasoner more flexibility to evolve.
5. Experiments
@@
In this section, we first introduce the experiment settings and training details of our method. We then present the main results of Meta-Reasoner with analysis regarding efficiency, rewards accumulation, and qualitative assessment of meta-reasoner output.
5.1. Experiments Setup
@@
We mainly focus on tasks that demand complex reasoning and often involve lengthy thinking processes for the correct solutions. These includes (1) 24-point game (Yao et al., 2023); (2) college-level scientific problem from SciBench (Wang et al., 2023) and (3) math questions based on theorems from TheoremQA (Chen et al., 2023). For SciBench, we focus only on the math-related subsets (i.e., diff, stat, and calc). Detailed explanations for each subset can be found in Wang et al. (2023). For TheormQA, we only consider the math subset that involves logical reasoning.
@@
We collect the training data for each task: (1) for the 24-point game, we random sample 50 queries from 4nums.com specifically excluding problems in ranks 901-1000 which were reserved for testing; (2) for TheoremQA, we randomly sample 30 mathematical reasoning queries from the dataset; (3) for SciBench, we randomly sample 30 queries from differential subsets including diff, stat, and calc from the entire dataset. These samples were used to iteratively update the LinUCB parameters for both fixed (KK=3 or KK=5 strategies) and dynamic strategy settings. We configure the training process using deterministic generation (nn=1, Top_k=1, temperature=0) with specific max_token limits for CoT generation (512), meta-reasoner feedback (256), progress reports (512), and reward model outputs (4). The LinUCB exploration parameter was set to cc=0.2. Experiments were ran for a maximum of TT=30 iterations for the 24-point game, scibench, and TT=100 for TheoremQA, with MAB parameters updated after each iteration based on a reward function that weighted objective completion (40%), progress quality (30%), efficiency (15%), and strategy alignment (15%). The reward function prompt is detailed in Figure 6-9.
Baselines.
We consider several established prompting methods as baselines as follows: (1) Chain-of-thought (CoT) (Wei et al., 2022): A prompting technique that encourages models to generate intermediate reasoning steps to enhance problem-solving capabilities. (2) Self-Consistent Chain of Thought (SC-CoT) (Wang et al., 2022): An extension of CoT that improves reasoning consistency by generating multiple reasoning chains and selecting the most consistent answer. (3) Multi-Chain Reasoning (MCR) (Yoran et al., 2023): enhances SC-CoT by having another LLM to assess and integrate content among the sampled reasoning chains to generate the final consistent answer. (4) Tree of Thoughts (ToT) (Yao et al., 2023): A method that explores multiple reasoning paths in a tree structure, allowing the model to consider various possibilities before arriving at a conclusion by tree search algorithms. (5) Reflexion (Shinn et al., 2023): A framework that enables models to reflect on their reasoning process, iteratively refining their answers based on feedback. (6) MACM (Lei et al., 2024): A multi-agent system to refine the reasoning based on iterative condition mining.
Backbone Models.
We consider both LLMs and the recent Large Reasoning Models (LRMs) for our experiments. For the LLMs, we consider the closed-source models like gpt-4o, gpt-4o-mini (between Nov 2025 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-3.1-8B-instruct from Meta, qwen-3-8B from Alibaba, phi-4 from Microsoft, gemini-experimental-1206 from Google, ds-r1-distill-llama-8B, ds-r1-distill-qwen-7B from Deepseek. For the LRMs, we consider the closed-source models like o1, o1-mini (In case we cannot break down the generation of o1 models through APIs, we cannot properly inject our meta-reasoner with o1-series models; we only provide the IO results for references). For the feature extraction mentioned in §4.3, we use text-embedding-3-small from OpenAI as the embedding model. To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models. We use the API service from OpenAI [^1] https://openai.com/ and OpenRouter [^2] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
5.2. Main Results
We compare the accuracy of different prompting methods across different backbone models on SciBench (as shown in Table 4), 24-points game (as shown in Table 2) and TheoremQA (as shown in Table 3). We found that basic prompting strategies, such as CoT and SC-CoT, show limited effectiveness, achieving only 4% and 9% accuracy on 24-point games, respectively. Incorporating IO strategy with “Best of 100” samples improves accuracy to 33%, but it remains far behind advanced methods. Strategies like ToT illustrate the importance of exploring broader reasoning paths, with accuracy increasing from 45% to 74% as the search breadth expands from 1 to 5. Advanced iterative methods, such as Reflexion (53%) and MACM (80%), further demonstrate the value of refined reasoning frameworks. Our proposed Meta-Reasoner outperforms these approaches, achieving 89% accuracy with GPT-4o-mini and 92% with GPT-4o, showcasing its ability to dynamically guide reasoning, correct errors, and focus resources effectively. Compared to specialized models like o1-mini, our method equipped with much cheaper and generalized models like GPT-4o-mini delivers comparable performance, demonstrating its adaptability and scalability.
Table 2. Performance of different models on 24-points Game (Yao et al., 2023) ($b$: search breadth)
Method Accuracy (%)
GPT-4o-mini + CoT (Yao et al., 2023) 4
GPT-4o-mini + SC-CoT (Yao et al., 2023) 9
GPT-4o-mini + IO (best of 100) (Yao et al., 2023) 33
GPT-4o-mini + CoT (best of 100) (Yao et al., 2023) 49
Gemini-Exp-1206 + IO (best of 100) (Yao et al., 2023) 38
Gemini-Exp-1206 + CoT (best of 100) (Yao et al., 2023) 60
GPT-4o-mini + ToT ($b=1$) (Yao et al., 2023) 32
GPT-4o-mini+ ToT ($b=5$) (Yao et al., 2023) 65
GPT-4o-mini + Reflexion (Shinn et al., 2023) 53
GPT-4o-mini + MACM (Lei et al., 2024) 80
GPT-4o-mini + Meta-Reasoner (our work) 89
GPT-4o + Meta-Reasoner (our work) 92
Gemini-Exp-1206 + Meta-Reasoner (our work) 94
o1-mini + IO 89
o1-preview + IO 93
Table 3. Performance of different models on TheoremQA (Chen et al., 2023)
Method Accuracy (%)
GPT-4o-mini + CoT 39.46
Gemini-Exp-1206 + CoT 43.12
GPT-4o-mini + Reflexion (Shinn et al., 2023) 74.32
GPT-4 Turbo + MACM (Lei et al., 2024) 79.41
GPT-4o-mini + Meta-Reasoner (our work) 84.13
Gemini-Exp-1206 + Meta-Reasoner (our work) 86.32
Overall, the Meta-Reasoner framework provides a compatible approach to improving reasoning-intensive tasks, combining high accuracy with dynamic and efficient problem-solving strategies. The results on SciBench and TheoremQA also demonstrate similar findings and show that Meta-Reasoner generally achieves better performance compared to the baselines and the results are consistent across different models.
Table 4. Performance of different models on SciBench Math Subsets (Wang et al., 2023)
@@ 5.2. Main Results
GPT-4o-mini + MACM (Lei et al., 2024) 54.78 67.13 65.77
GPT-4o + MACM (Lei et al., 2024) 61.42 78.32 76.72
@@
5.3. Ablation Study
We further conduct an ablation study to analyze each component contribution of Meta-Reasoner. In specific, we consider the following setup: (1) w/o progress report: we replace the progress reporting process with directly considering the entire CoT history without summarization; (2) w/o MAB: instead of using MAB to select the proper strategy, we directly leverage an LLM to the decision making to provide the proper strategy for LRM reasoning.
Table 5. Ablation study of Meta-Reasoner. Direct arm selection refers to prompting LLM to directly select a strategy based on recent progress report.
Model Variant Game-of-24 TheoremQA
@@ 5.3. Ablation Study
In Table 5, we show that when removing progress reporting (“w/o Progress Report”), the overall performance moderately degrades and we hypothesize it is due to the concise intermediate summarizations can help the Meta-reasoner only consider the high-level strategy instead of being confused with too much details of the reasoning process. We also find that removing the MAB brings a more pronounced effect, especially when strategy selection falls back to a direct chain-of-thought approach (“w/o MAB (CoT)”). It verifies the effect of our meta-reasoner module to help the model stay on track for getting an optimal solution. In Table 6, we compare fixed and dynamic bandit variants on the game of 24 and theoremQA. We find that using a fixed set of strategies (e.g., $K=3$ and $K=5$) yields lower performance compared to the dynamic approach which adaptively explores more strategies (shown by larger unique strategies). The results highlight the benefit of flexibly allocating diverse reasoning strategies using LLM in-context learning capabilities.
Figure 3. Cumulative reward of different settings across iteration. We compare our method using LinUCB with baseline (direct arm selection), and random search methods across two tasks—Game of 24 (top row) and TheoremQA (bottom row) using GPT-4o-mini (left) and Gemini-Exp-1206 (right).
(a) Game of 24 Task.
(b) TheoremQA Task
Figure 4. Inference time heatmap across different methods.
5.4. Analysis
5.4.1. Effectiveness of Meta-reasoner.
Figure 3 demonstrates the cumulative rewards across iterations. We compare our MAB-based approach with a baseline that directly prompts an LLM to select an arm (or “strategy”), referred to as Baseline (Direct Arm Selection); the prompt details are in Figures 69. Results show that the MAB-based meta-reasoner (using LinUCB (Li et al., 2010)) consistently outperforms both direct LLM decision-making and random search across two tasks (Game of 24 and TheoremQA) and two model scales (GPT-4o-mini and Gemini-Exp-1206). While direct LLM prompting yields reasonable initial performance and random search requires minimal setup, neither approach effectively balances exploration and exploitation. In contrast, the MAB updating strategy leverages feedback from prior iterations to adaptively refine action selection (e.g., choosing an appropriate strategy based on CoT reasoning), steadily increasing cumulative rewards.
Table 6. Fixed vs. Dynamic Bandit Variants over GPT-4o-mini. #US: Number of Unique Strategies.